Use mathematical induction to prove the following.
N^3 < or = (N+1)^2 ; N> or = 2.

Answer by ikleyn(52794)

.
Use mathematical induction to prove the following.
N^3 < or = (N+1)^2 ; N> or = 2.
~~~~~~~~~~~~~~~~~~~



The problem' statement is  FATALLY  WRONG.

        A counter-example is  N = 3:   N^3 = 27;   (N+1)^2 = 4^2 = 16.


Do not post  FALSE  STATEMENTS  to prove.